PLEASE HELP ME SOLVE THIS EQUATION 3K-5=7K-21
Answer by tutorcecilia(2144)

3K-5=7K-21
3K-3K-5=7K-3K-21
-5=4K-21
-5+21=4K-21+21
16=4K
16/4=4K/4
4=K
.
check by plugging (K=4) back into the original equation.